OpenStudy (anonymous):

Please confirm!!!! Will Medal (and must be correct) F(x) = 7 - 2x = All real numbers greater than 7 (???)

OpenStudy (hermom1967):

Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 7−2x=F(x) Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides. −2x=−7+F(x) Reorder the polynomial −7+F(x) alphabetically from left to right, starting with the highest order term. −2x=F(x)−7

OpenStudy (anonymous):

D:

OpenStudy (hermom1967):

Divide each term in the equation by −2. \[-\frac{ 2x }{ -2 }=\frac{ F(x) }{ -2 }-\frac{ 7 }{ -2 }\]

OpenStudy (hermom1967):

Move the minus sign from the denominator to the front of the expression. \[\frac{ 2x }{ 2 }=\frac{ F(x) }{ -2 }-\frac{ 7 }{ -2 }\]

OpenStudy (anonymous):

So less than 5?

OpenStudy (hermom1967):

Cancel the common factor of 2 the expression \[\frac{ 2x }{ 2 }\] \[x=\frac{ F(x) }{ -2 }-\frac{ 7 }{ -2 }\]

OpenStudy (anonymous):

All real numbers less than 5?

OpenStudy (hermom1967):

No

OpenStudy (anonymous):

Less than seven then...?

OpenStudy (anonymous):

A. All real numbers less than 5 B. All real numbers less than 7 C. All real numbers greater than 7 D. All real numbers greater than 5

OpenStudy (hermom1967):

Move the minus sign from the denominator to the front of the expression. \[x=-\frac{ F(x) }{ 2 }-\frac{ 7 }{ -2 }\] Move the minus sign from the denominator to the front of the expression. \[x=-\frac{ F(x) }{ 2 }+\frac{ 7 }{ 2 }\] Reorder the polynomial alphabetically from left to right, starting with the highest order term. \[x=-\frac{ 7 }{ 2 }-\frac{ F(x) }{ 2 }\]

OpenStudy (hermom1967):

Are you looking for the range?

OpenStudy (anonymous):

Thanks. Chose an answer and moved on.